Grade 11Physical Chemistry

In a face centred cubic lattice, atom A occupies the corner positions and atom B occupies the face centre positions. If one atom of B is missing from one of the face centred points, the formula of the compound is : (1) 2A B (2) 23A B (3) 25A B (

tanisha singh
12 Years agoGrade 11

2 Answers

Sunil Kumar FP
11 Years ago
As the molecule is having FCC structure,therefore
let the formula of the compound be AxBy
no of atom of A=1/8*8=1
number of atom of B=6*1/2 -1/2=5/2,As one atom is missing and ,where the contribution of each atom in FCC is ½.
Thus the formula is A1B5/2 or A2B5
Kushagra Madhukar
5 Years ago
Dear student,
Please find the solution to your problem below.
 
Let the formula of the compound be AxBy
no of atoms of A = 1/8*8 = 1
number of atom of B = 6*1/2 –1/2=5/2
 
Hence, The Molecular formula is A1B5/2 or A2B5
 
Thanks and regards,
Kushagra